# GEORCE: A Fast New Control Algorithm for Computing Geodesics

### v1

@@ Abstract
Computing geodesics for Riemannian manifolds is a difficult task that often relies on numerical approximations. However, these approximations tend to be either numerically unstable, have slow convergence, or scale poorly with manifold dimension and number of grid points. We introduce a new algorithm called GEORCE that computes geodesics via a transformation into a discrete control problem. We show that GEORCE has global convergence and quadratic local convergence. In addition, we show that it extends to Finsler manifolds. For both Finslerian and Riemannian manifolds, we thoroughly benchmark GEORCE against several alternative optimization algorithms and show empirically that it has a much faster and more accurate performance for a variety of manifolds, including key manifolds from information theory and manifolds that are learned using generative models.
@@ 1 Introduction
Geodesics, or locally shortest paths, are the key computational tools for engaging with data residing on a Riemannian manifold. The lengths of geodesic curves inform us about manifold distances. Likewise, the exponential and logarithm maps, which generalize addition and subtraction, are rooted in geodesics [22]. Unfortunately, the solution to the boundary-value problem for geodesics are only available in closed form for the simplest manifolds, and numerical approximations are therefore required in practice.
Such approximations are typically achieved by minimizing the energy functional over a suitable set of curves connecting the given boundary points and parametrized by $\gamma:[0,1]\rightarrow\mathcal{M}$ [23]
@@
where $G$ denotes the metric matrix function of the manifold in local coordinates. This energy is commonly minimized with general-purpose gradient-based optimization methods [24, 25, 12] or, in particular when considering the initial value problem and using shooting methods, by solving the system of ordinary differential equations that arise by applying the Euler-Lagrange equations [26, 27]. Unfortunately, both approaches are often numerically brittle, exhibit slow convergence, or scale poorly with the number of grid points and manifold dimensions.
Figure 1: Comparison between GEORCE and baseline methods for computing locally length minimizing geodesics between given point pairs on four different manifolds. All algorithms are terminated if the $\ell^{2}$-norm of the gradient of the discretized energy functional (5) is less than $10^{-4}$, or the number of iterations exceeds $100$. Note that sometimes the respective solutions overlap, making it difficult to distinguish them. Also note, that in the case of the torus the initial curve determines the direction around the hole so that the obtained geodesics follow this direction and are therefore clearly not the globally shortest geodesics between the given start and end point. Experimental details and benchmark data are in Appendix E.
@@ 2 Background and related work
A Riemannian manifold, $(\mathcal{M},g)$, can indirectly be considered as a smooth $n$-dimensional surface in a Euclidean space of sufficiently high dimension, see [28, 29]. In this way $\mathcal{M}$ inherits its metric $g$ from the ambient space. All the intrinsic geometry of $(\mathcal{M},g)$, including curvature and geodesic distances, are encoded in the metric $g$. The metric can be – and often is – constructed independently of the formal embedding alluded to above, namely directly as a symmetric positive definite matrix valued function on the domain of a parametrization of the manifold in question. The metric, $g:T_{x}\mathcal{M}\times T_{x}\mathcal{M}\rightarrow\mathbb{R}$, defines an inner product that smoothly varies over the tangent spaces of $\mathcal{M}$ such that in a local chart $g(v,w)=v^{\top}G(x)w$, where $G$ is the metric matrix function evaluated in $x\in\mathcal{M}$ for $v,w\in T_{x}\mathcal{M}$. The tangent space, $T_{x}\mathcal{M}$, consists of tangents to all curves at $x\in\mathcal{M}$ [23].
A geodesic can then be defined using the inner product in each tangent space. The length of a curve, $\gamma:U\subseteq[0,1]\rightarrow\mathcal{M}$ parameterized by $t\in[0,1]$ is given by
@@
with the corresponding energy given by eq. 1. A geodesic is a smooth curve that locally minimizes the length between two points $x,y\in\mathcal{M}$ and gives rise to the Riemannian distance, $\mathrm{dist}:\mathcal{M}\times\mathcal{M}\rightarrow\mathbb{R}_{+}$, which is defined as the length of the corresponding shortest geodesic between the given two points on the manifold [23]. It can be shown that if a curve $\gamma$ locally minimizes the energy (1), then the curve has constant velocity and locally minimizes the length [30]. For general Riemannian manifolds, existence and uniqueness are only guaranteed in a star-shaped neighborhood around any point $x\in\mathcal{M}$ [23]. All points, where the geodesics from a point $x\in\mathcal{M}$ are no longer length minimizing, are denoted the cut locus of $x$ [23]. In this paper, we assume that the manifold is geodesically complete such that at least one length-minimizing geodesic exists between any two points $x,y\in\mathcal{M}$.
A Riemannian manifold can be equipped with a unique metrically compatible connection, $\nabla$, known as the Levi-Civita Connection [23], where uniqueness and existence are guaranteed by completeness of the metric [23]. An equivalent way to define geodesics is to define these as curves, $\gamma$, with zero acceleration using the Levi-Civita connection as the covariant derivative, i.e., $\nabla_{\dot{\gamma}(t)}\dot{\gamma}(t)=0$. In a local chart, this gives rise to the following (usually non-linear) ode-system [23]
@@
written in Einstein notation [^1] In Einstein notation, upper and lower indices constitute a sum, i.e. $a^{i}b_{i}=\sum_{i}a_{i}b_{i}$, where $\left\{\Gamma_{ij}^{k}\right\}_{i,j,k}$ denote the Christoffel symbols. The exponential map $\mathrm{Exp}_{x}(v):T_{x}\mathcal{M}\rightarrow\mathcal{M}$ is then defined as $\mathrm{Exp}_{x}(v)=\gamma(1)$, where $\gamma$ is a geodesic with $\gamma(0)=x\in\mathcal{M}$ and initial velocity vector, $v\in T_{x}\mathcal{M}$, i.e. $\dot{\gamma}(0)=v$. Within the cut locus of $x\in\mathcal{M}$ the exponential map, $\mathrm{Exp}_{x}$, is a diffeomorphism and its inverse is the logarithmic map, $\mathrm{Log}_{x}:\mathcal{M}\rightarrow T_{x}\mathcal{M}$.
@@
The energy functional in eq. 1 can be discretized and minimized using standard optimization methods. Consider fixed start and end points $x_{0}=a\in\mathcal{M}$ and $x_{T}=b\in\mathcal{M}$, and let $x_{0:T}:=\{x_{t}\}_{t=0}^{T}$ denote a discretized version of the candidate curve $\gamma$ in $T+1$ points. Minimizing the energy functional can be re-written as a discrete constrained optimization problem, i.e., subdivision of $[0,1]$ into $T$ subintervals, in the chosen local coordinate system [25, 12]
@@
where $G\left(\cdot\right)$ denotes the metric matrix function in local coordinates, and the tangent vectors of the coordinate expression for $\gamma$ are discretized as $\dot{\gamma}(t)\approx\left(x_{t+1}-x_{t}\right)T$. Note that the constant scaling by $\nicefrac{{1}}{{2}}$ is omitted since it does not affect the optimal solution. With this formulation, the curve is approximated as a piece-wise linear function. This is not restrictive since $\gamma$ can be approximated into a curve for any given precision $T$ or converted to a smooth function, e.g., using splines [24].
The optimization problem in eq. 5 can be solved using standard solvers in the variables $x_{1:(T-1)}$. Gradient descent converges to a local minimum (global convergence), but will have local linear convergence, while a Newton method or quasi-Newton method will have local quadratic and super-linear convergence, respectively, but will involve solving a linear equation system and will, in general, not have global convergence [33].
@@
Figure 2: The estimates of bvp-solvers with different integration methods compared to GEORCE on $\mathbb{S}^{n}$ for $n=2,3,5,10$. The bvp-solvers minimize the squared error to the boundary condition using BFGS [37, 38, 39, 40]. All methods are terminated if they take more than 24 hours or use more than 10 GB of memory on a CPU. Details on the experiment is found in Appendix E.1.
Directly minimizing the energy in Eq. 5 can, e.g., be done using off-the-self solvers like ADAM [41] or BFGS [37, 38, 39, 40]. However, these methods tend to have slow convergence and scale poorly in the manifold dimension and number of grid points (see Sec. 5). Alternatively, the geodesic can be estimated by discretizing the smooth manifold by simplicial complexes [42], for which there exist different geodesic solvers [43, 44, 45, 46, 47]. Although discretized manifolds appear naturally in computational geometry, the approach is not scalable to discretize high-dimensional smooth manifolds.
@@ 3 GEORCE
To develop an efficient primal-dual algorithm for computing geodesics, we will first reformulate the discretized energy functional (5) as a discrete control problem. We then derive the necessary conditions for a minimum for the control problem using techniques from optimal control to decompose the discrete control problem into strictly convex subproblems; from this, we deduce an iterative scheme to compute geodesics. Finally, we prove that our proposed algorithm exhibits global convergence and local quadratic convergence.
As a starting point, we reformulate the problem of finding a minimum of the discrete energy functional (5) to a control problem
@@
See Appendix A.2. ∎
Proposition 1 decomposes the minimization problem in Eq. 6 into a system of equations for each time step $t$. Note that these equations are necessary conditions for the optimum for eq. 6 and are sufficient in the case where the objective function in eq. 6 is convex in both $x_{t}$ and $u_{t}$. However, this is generally not the case.
@@
$\begin{split}&2G_{t}u_{t}+\mu_{t}=0,\quad t=0,\dots,T-1,\\ &\nu_{t}+\mu_{t}=\mu_{t-1},\quad t=1,\dots,T-1,\\ &\sum_{t=0}^{T-1}u_{t}=b-a,\\ \end{split}$ (9)
@@
See Appendix A.3. ∎
@@
4:while $\ \textbf{do}\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big{|}}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$
@@
8:     $u_{t}\leftarrow-\frac{1}{2}G_{t}^{-1}\left(\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j}\right)$ for $t=0,\dots,T-1$
9:     $x_{t+1}\leftarrow x_{t}+u_{t}$ for $t=0,\dots,T-1$
10:     Using line search find $\alpha^{*}$ for the following optimization problem with the discrete energy functional $E$
$\begin{split}\alpha^{*}=\operatorname*{arg\,min}_{\alpha}\quad&E\left(x_{0:T}\right)\quad\text{(exact line search)}\\ \text{s.t.}\quad&x_{t+1}=x_{t}+\alpha u_{t}+(1-\alpha)u_{t}^{(i)},\quad t=0,\dots,T-1,\\ &x_{0}=a.\end{split}$
11:     Set $u_{t}^{(i+1)}\leftarrow\alpha^{*}u_{t}+(1-\alpha^{*})u_{t}^{(i)}$ for $t=0,\dots,T-1$
12:     Set $x_{t+1}^{(i+1)}\leftarrow x_{t}^{(i+1)}+u_{t}^{(i+1)}$ for $t=0,\dots,T-1$
13:     $i\leftarrow i+1$
14:end while
15:return $x_{t}$ for $t=0,\dots,T-1$
Proposition 2 immediately leads to an iterative updating of $\left(x_{t}^{(i)},u_{t}^{(i)}\right)$ to estimate the geodesic, which we denote the GEORCE-algorithm. GEORCE is described in pseudo-code in Algorithm 1.
@@
We find that an appropriate step size is necessary to ensure global convergence for GEORCE (See Proposition 3), and therefore we update $u_{t}^{(i+1)}\leftarrow\alpha u_{t}^{(i+1)}+(1-\alpha)u_{t}^{(i)}$ in iteration $i$. Here, $\alpha$ denotes the step size, which we estimate with line search. One approach is to estimate the optimal step size in each iteration, but, in our experience, this is unnecessarily expensive. We instead use soft line search, which merely ensures that the objective function decreases in each iteration [33]. We implement this using backtracking, where an initial step size is multiplied by a decay rate, $10<\rho<$ 1, until some condition is met, e.g., the Wolfe condition [48], the Armijo condition [49], or the curvature condition [48]. In practice, we find that the Armijo condition [49] with $\rho=0.5$ is sufficient for most manifolds, although the decay rate can be fine-tuned to increase the performance of GEORCE.
@@
Convergence results
In the following, we show that GEORCE is guaranteed to converge to a local minimum, i.e., it exhibits global convergence.
Proposition 3 (Global convergence).
Let $E^{(i)}$ be the value of the discretized energy functional for the solution after iteration $i$ (with line search) in GEORCE. If the starting point $\left(x^{(0)},u^{(0)}\right)$ is feasible in the sense that the geodesic candidate has the correct start and end points, then the series $\left\{E^{(i)}\right\}_{i>0}$ will converge to a (local) minimum.
Proof.
See Appendix A.5 for the proof. ∎
In theory, GEORCE could jump between two local minima with the same objective value unless the local minimum is strongly unique, as described in Appendix A.1, and seen in conjugate points [23]. Proposition 3 ensures that GEORCE converges to a local minimum (global convergence) similar to gradient descent. In contrast, gradient descent has only linear local convergence, but we show in the following that under sufficient regularity of the discretized energy functional (see Appendix A.1), GEORCE has quadratic local convergence.
Proposition 4 (Local quadratic convergence).
Under sufficient regularity of the discretized energy functional (see Appendix A.1) it holds that if the discretized energy functional has a strongly unique (local) minimum point $z^{*}$ (see Appendix A.1) and locally the step-size $\alpha^{*}=1$, i.e., no line search, then GEORCE has locally quadratic convergence, i.e.,
$\exists\epsilon>0:\quad\exists c>0:\quad\forall z^{(i)}\in B_{\epsilon}(z^{*}):\quad\left\lVert z^{(i+1)}-z^{*})\right\rVert\leq c\left\lVert z^{(i)}-z^{*}\right\rVert^{2},$
where $z^{(i)}=(x^{(i)},u^{(i)})$ is the solution from GEORCE in iteration $i$, and $B_{\epsilon}\left(z^{*}\right)$ is the ball with radius $\epsilon>0$ and center $z^{*}$.
Proof.
See Appendix A.6 for the proof and Appendix A.1 for the assumptions. ∎
Fig. 3 demonstrates the convergence by plotting the $\ell^{2}$-norm of the gradient of the discretized energy functional in each iteration for four different manifolds. Note that by formulating the geodesic boundary problem as a control problem, we estimate the discretized tangent vectors along the curve $\{u_{t}\}_{t=0}^{T}$. Modulo scaling $u_{0}$ will in the limit of a sufficiently fine grid correspond to the logarithmic map, $\mathrm{Log}_{a}b$.
@@
Table 1: Convergence, complexity, and use of higher order derivatives for different optimization algorithms in each iteration, where $T$ is the number of grid points and $d$ is the manifold dimension.

### v2

@@ Abstract
Computing geodesics for Riemannian manifolds is a difficult task that often relies on numerical approximations. However, these approximations tend to be either numerically unstable, have slow convergence, or scale poorly with manifold dimension and number of grid points. We introduce a new algorithm called GEORCE that computes geodesics in a local chart via a transformation into a discrete control problem. We show that GEORCE has global convergence and quadratic local convergence. In addition, we show that it extends to Finsler manifolds. For both Finslerian and Riemannian manifolds, we thoroughly benchmark GEORCE against several alternative optimization algorithms and show empirically that it has a much faster and more accurate performance for a variety of manifolds, including key manifolds from information theory and manifolds that are learned using generative models.
@@ 1 Introduction
Geodesics, or locally shortest paths, are the key computational tools for engaging with data residing on a Riemannian manifold. The lengths of geodesic curves inform us about manifold distances. Likewise, the exponential and logarithm maps, which generalize vector addition and subtraction, are rooted in geodesics [22, Chapter 9]. Unfortunately, the solution to the boundary-value problem for geodesics are only available in closed form for the simplest manifolds, and numerical approximations are therefore required in practice.
Such approximations are typically achieved by minimizing the energy functional over a suitable set of curves connecting the given boundary points and parametrized by $\gamma:[0,1]\rightarrow U$ [23, Page 194]
@@
where $G$ denotes the metric matrix function of the manifold in local coordinates in a chart $\phi:U\rightarrow\mathcal{M}$, where $U$ is an open subset of $\mathbb{R}^{d}$. This energy is commonly minimized with general-purpose gradient-based optimization methods [24, 11, 12] or, in particular when considering the initial value problem and using shooting methods, by solving the system of ordinary differential equations that arise by applying the Euler-Lagrange equations [25, 26]. Unfortunately, both approaches are often numerically brittle, exhibit slow convergence, or scale poorly with the number of grid points and manifold dimensions.
Figure 1: Comparison between GEORCE and baseline methods for computing locally length minimizing geodesics between given point pairs on four different manifolds. All algorithms are terminated if the $\ell^{2}$-norm of the gradient of the discretized energy functional (5) is less than $10^{-4}$, or the number of iterations exceeds $100$. Note that sometimes the respective solutions overlap, making it difficult to distinguish them. Also note, that in the case of the torus the initial curve determines the direction around the hole so that the obtained geodesics follow this direction and are therefore clearly not the globally shortest geodesics between the given start and end point. Experimental details and benchmark data are in Appendix F.
@@ 2 Background and related work
A Riemannian manifold, $(\mathcal{M},g)$, can indirectly be considered as a smooth $n$-dimensional surface in a Euclidean space of sufficiently high dimension, see [27, 28]. In this way $\mathcal{M}$ inherits its metric $g$ from the ambient space. All the intrinsic geometry of $(\mathcal{M},g)$, including curvature and geodesic distances, are encoded in the metric $g$. The metric can be – and often is – constructed independently of the formal embedding alluded to above, namely directly as a symmetric positive definite matrix valued function on the domain of a parametrization of the manifold in question. The metric, $g:T_{x}\mathcal{M}\times T_{x}\mathcal{M}\rightarrow\mathbb{R}$, defines an inner product that smoothly varies over the tangent spaces of $\mathcal{M}$ such that in a local chart $g(v,w)=v^{\top}G(x)w$, where $G$ is the metric matrix function evaluated in $x\in\mathcal{M}$ for $v,w\in T_{x}\mathcal{M}$. The tangent space, $T_{x}\mathcal{M}$, consists of tangents to all curves at $x\in\mathcal{M}$ [23, Page 7 Definition 2.6].
A geodesic can then be defined using the inner product in each tangent space. The length of a curve, $\gamma:[0,1]\rightarrow U$ parameterized by $t\in[0,1]$ is given by
@@
where $\phi:U\rightarrow\mathcal{M}$ is a local chart for an open set $U\subset\mathbb{R}^{d}$. A Riemannian manifold can be equipped with a unique metrically compatible connection, $\nabla$, known as the Levi-Civita Connection [23, Page 55 Theorem 3.6], where uniqueness and existence are guaranteed by completeness of the metric [23, Page 63 Lemma 2.3]. Note that curves are geodesics if and only if they are critical points of the energy functional, i.e., the first variation of the energy functional is zero for all variations with fixed endpoints assuming that the end points are not conjugate points [23, Page 196 Proposition 2.5]. This also holds in the Finslerian case [29, Page 81].
An equivalent way to define geodesics is to define these as curves, $\gamma$, with zero acceleration using the Levi-Civita connection as the covariant derivative, i.e., $\nabla_{\dot{\gamma}(t)}\dot{\gamma}(t)=0$. In a local chart, this gives rise to the following (usually non-linear) ode-system [23, Page 62]
@@
written in Einstein notation [^1] In Einstein notation, upper and lower indices constitute a sum, i.e. $a^{i}b_{i}=\sum_{i}a_{i}b_{i}$, where $\left\{\Gamma_{ij}^{k}\right\}_{i,j,k}$ denote the Christoffel symbols. The exponential map $\mathrm{Exp}_{x}(v):T_{x}\mathcal{M}\rightarrow\mathcal{M}$ is then defined as $\mathrm{Exp}_{x}(v)=\gamma(1)$, where $\gamma$ is a geodesic with $\gamma(0)=x\in\mathcal{M}$ and initial velocity vector, $v\in T_{x}\mathcal{M}$, i.e. $\dot{\gamma}(0)=v$. Within the injectivity radius of $x$ the exponential map, $\mathrm{Exp}_{x}$, is a diffeomorphism and its inverse is the logarithmic map, $\mathrm{Log}_{x}:\mathcal{M}\rightarrow T_{x}\mathcal{M}$.
@@
The energy functional in Eq. 1 can be discretized and minimized using standard optimization methods. Consider fixed start and end points $x_{0}=a\in U$ and $x_{T}=b\in U$, and let $x_{0:T}:=\{x_{t}\}_{t=0}^{T}$ denote a discretized version of the candidate curve $\gamma$ in $T+1$ points in a local chart $U$. Minimizing the energy functional can be re-written as a discrete constrained optimization problem, i.e., subdivision of $[0,1]$ into $T$ subintervals, in the chosen local chart [11, 12]
@@
where $G\left(\cdot\right)$ denotes the metric matrix function in local coordinates, and the tangent vectors of the coordinate expression for $\gamma$ are discretized as $\dot{\gamma}(t)\approx\left(x_{t+1}-x_{t}\right)T$. Note that the constant scaling by $\nicefrac{{1}}{{2}}$ is omitted since it does not affect the optimal solution. With this formulation, the curve is approximated as a piece-wise linear function. This is not restrictive since $\gamma$ can be approximated into a curve for any given precision $T$ or converted to a smooth function, e.g., using splines [24]. Note that by approximating the functional energy by forward differences, the computational error is linear in $\nicefrac{{1}}{{T}}$, that is, in the form $\mathcal{O}\left(\nicefrac{{1}}{{T}}\right)$.
Figure 2: The estimates of bvp-solvers with different integration methods compared to GEORCE on $\mathbb{S}^{n}$ for $n=2,3,5,10$. The bvp-solvers minimize the squared error to the boundary condition using BFGS [32, 33, 34, 35]. All methods are terminated if they take more than 24 hours or use more than 10 GB of memory on a CPU. Details on the experiment is found in Appendix F.1.
Similarly to the energy functional in Eq. 1, it is assumed that a stationary point to the discretized energy functional in Eq. 5 is either a local or global minimum point. Thus, the number of points, T, in Eq. 5 needs to be sufficiently high to achieve this. If this is not fulfilled, then there may exist saddle points for the discretized energy functional.
The optimization problem in Eq. 5 can be solved using standard solvers in the variables $x_{1:(T-1)}$. Gradient descent converges to a local minimum (global convergence), but will have local linear convergence, while a Newton method or quasi-Newton method will have local quadratic and super-linear convergence, respectively, but will involve solving a linear equation system and will, in general, not have global convergence [36, Chapter 8].
@@
Directly minimizing the energy in Eq. 5 can, e.g., be done using off-the-shelf solvers like ADAM [40] or BFGS [32, 33, 34, 35]. However, these methods tend to have slow convergence and scale poorly in the manifold dimension and number of grid points (see Sec. 5). Alternatively, the geodesic can be estimated by discretizing the smooth manifold by simplicial complexes [41], for which there exist different geodesic solvers [42, 43, 44, 45, 46]. Although discretized manifolds appear naturally in computational geometry, the approach is not scalable to discretize high-dimensional smooth manifolds.
@@ 3 GEORCE
To develop an efficient primal-dual algorithm for computing geodesics, we will first reformulate the discretized energy functional (5) as a discrete control problem in a local chart. We then derive the necessary conditions for a minimum for the control problem using techniques from optimal control to decompose the discrete control problem into strictly convex subproblems; from this, we deduce an iterative scheme to compute geodesics. Finally, we prove that our proposed algorithm exhibits global convergence and local quadratic convergence.
We will assume the following for our algorithm
  • We assume access to a local chart of the manifold and assume that the entire candidate curve of the geodesic can be represented within the same chart.
  • All computations are performed within the local coordinate chart $U$, which is defined based on the induced metric of the Riemannian manifold.
  • For computational efficiency, we will leverage the Euclidean properties of the local chart (i.e., straight-line interpolation within $U$).
  • We will assume that the end points are not conjugate points, and that the discretization is sufficiently fine such that a critical point of the discretized energy functional in Eq. 5 is either a local or global minimum point.
3.1 Deriviaton of GEORCE
As a starting point, we reformulate the problem of finding a minimum of the discrete energy functional (5) to a control problem in a local chart
@@ 3.1 Deriviaton of GEORCE
The Hamiltonian of Eq. 6 is
$H_{t}(x_{t},u_{t},\mu_{t})=u_{t}^{\top}G(x_{t})u_{t}+\mu_{t}^{\top}(x_{t}+u_{t}).$
Proposition 1 decomposes the minimization problem in Eq. 6 into a system of equations for each time step $t$. Note that these equations are necessary conditions for the optimum for Eq. 6 and are sufficient in the case where the objective function in Eq. 6 is convex in both $x_{t}$ and $u_{t}$ [48, Page 244]. However, this is generally not the case, but due to the assumption applying a sufficiently fine discretization the solution to Eq. 7 will be a local or global minimum point.
@@
[…]
@@
From the update formula $\nu_{t}+\mu_{t}=\mu_{t-1}$ for $t=1,\dots,T-1$ we deduce that $\mu_{t}=\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j}$. From the update scheme $2G_{t}u_{t}+\mu_{t}=0$ we deduce that $u_{t}=-\frac{1}{2}G_{t}^{-1}\mu_{t}$, since $G_{t}$ is positive definite and hence has a well-defined inverse. Noting that $\sum_{t=0}^{T-1}u_{t}=b-a$ we can re-write the iterative formulas as
which proves the result. ∎
@@
4:while $\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big|}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$ do
@@
8:  $\tilde{u}_{t}\leftarrow-\frac{1}{2}G_{t}^{-1}\left(\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j}\right)$ for $t=0,\dots,T-1$
9:  Using line search find $\alpha^{*}$ for the following optimization problem with the discrete energy functional $E$
$\begin{split}\alpha^{*}=\operatorname*{arg\,min}_{\alpha}\quad&E\left(\tilde{x}_{0:T}\right)\quad\text{(exact line search)}\\ \text{s.t.}\quad&\tilde{x}_{t+1}=\tilde{x}_{t}+\alpha\tilde{u}_{t}+(1-\alpha)u_{t}^{(i)},\quad t=0,\dots,T-1,\\ &\tilde{x}_{0}=a.\end{split}$
10:  Set $u_{t}^{(i+1)}\leftarrow\alpha^{*}\tilde{u}_{t}+(1-\alpha^{*})u_{t}^{(i)}$ for $t=0,\dots,T-1$
11:  Set $x_{t+1}^{(i+1)}\leftarrow x_{t}^{(i+1)}+u_{t}^{(i+1)}$ for $t=0,\dots,T-1$
12:  $i\leftarrow i+1$
13:end while
14:return $x_{t}$ for $t=0,\dots,T-1$
[…]
As we show below, GEORCE exhibits global convergence, which means that the algorithm will converge to a local minimum. If there are multiple local minima, then GEORCE will not necessarily find the global minimum. To investigate whether GEORCE has found the global minimum, one can apply GEORCE for different initial curves to compute locally-length minimizing curves and choose the curve with the shortest length as the global minimum. In Appendix C, we illustrate the effect of different choices of initialization curves and how it affects the solution using GEORCE.
@@
We find that an appropriate step size is necessary to ensure global convergence for GEORCE (See Proposition 5), and therefore we update $u_{t}^{(i+1)}\leftarrow\alpha u_{t}^{(i+1)}+(1-\alpha)u_{t}^{(i)}$ in iteration $i$. Here, $\alpha$ denotes the step size, which we estimate with line search. One approach is to estimate the optimal step size in each iteration, but, in our experience, this is unnecessarily expensive. We instead use soft line search, which merely ensures that the objective function decreases in each iteration [36]. We implement this using backtracking, where an initial step size is multiplied by a decay rate, $0<\rho<1$, until some condition is met, e.g., the Wolfe condition [49], the Armijo condition [50], or the curvature condition [^2] Note that the term “curvature condition” is rather misleading in a Riemannian context as it refers to conditions on the derivative of the objective function and not the Riemannian curvature. [49]. In practice, we find that the Armijo condition [50] with $\rho=0.5$ is sufficient for most manifolds, although the decay rate can be fine-tuned to increase the performance of GEORCE.
@@
Table 1: Convergence, complexity, and use of higher order derivatives for different optimization algorithms in each iteration, where $T$ is the number of grid points and $d$ is the manifold dimension. A checkmark indicates that the algorithm exhibits the property, while a cross indicates that it is not the case.